OpenStudy (anonymous):

Find the indefinite integral:

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\frac{ e^{y} }{ 2-e^{y} } dy\]

OpenStudy (anonymous):

i know i'll be using substitution.

OpenStudy (turingtest):

Let\[u=2-e^y\]then\[du=?\]

OpenStudy (anonymous):

e^y y dy?

OpenStudy (turingtest):

\[u=2-e^y\implies du=-e^ydy\implies e^ydy=-du\]so what is the integral now?

OpenStudy (anonymous):

\[\int\limits_{?}^{?} 1/du du\]

OpenStudy (anonymous):

well minus?

OpenStudy (turingtest):

how did you get two du's ?

OpenStudy (anonymous):

\[\int\limits_{e^y}}^{2-e^{y}} e^{\]

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ e^{y} }{ 2-e{y} } e^{y}\]

OpenStudy (turingtest):

where is the other e^y coming from?

OpenStudy (anonymous):

thats my du.

OpenStudy (anonymous):

well dy.

OpenStudy (anonymous):

it should be negative.

OpenStudy (anonymous):

is the answer ln|2-e^y|+c?

OpenStudy (turingtest):

you dropped the negative sign, but aside from that the answer is correct